Heat flows through a wooden board 30 mm thick at 0.0086 W/cm³ ,when one of its sides is 20˚C warmer than the other. Determine the thermal conductivity of the wood.
1
Expert's answer
2020-05-24T18:01:12-0400

Thermal conductivity

"K = \frac {Q} {A\times\Delta T}"


"K = \frac {0.0086} {3\times1\times(20+273)}"


"K = 9.78 \times 10 ^{-6} \frac {W} {cm^{2}\times K}"


"K = 0.0978 \frac {W} {m^{2}\times K}"